OpenStudy (anonymous):

how does -4(sqrt3/2)= -2sqrt3

OpenStudy (anonymous):

\[-4(\sqrt{3}/2)= -2\sqrt{3}\]

OpenStudy (anonymous):

to break it down the equation on the left -4(sqrt3/2) = -2.2.(sqrt3/2). the 2s cancel each other out and u get the equation on the right.

OpenStudy (anonymous):

oh i see thanks